OpenStudy (anonymous):

1,440 is 120 percent of

OpenStudy (anonymous):

are you looking for answers, or explanations of how to do these?

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

you generally shouldn't just flood the questions forum with like 10 questions that are almost the same looking for answers. That's just dishonest.

OpenStudy (anonymous):

120/100 * x = 1440 x=1440*100/120=1200

OpenStudy (anonymous):

learn how to do it on one or two, then figure the rest out on your own.

OpenStudy (anonymous):

ohkayy @ hickknijaa